### v1

# Fulde-Ferrell-Larkin-Ovchinnikov States and Topological Bogoliubov Fermi Surfaces in Altermagnets: an Analytical Study

Abstract
We present an analytical study of the ground-state phase diagram for dilute two-dimensional spin-1/2 Fermi gases exhibiting $d$-wave altermagnetic spin splitting under $s$-wave pairing. Within the Bogoliubov-de Gennes mean-field framework, four distinct phases are identified: a Bardeen-Schrieffer-Cooper-type superfluid, a normal metallic phase, a nodal superfluid with topological Bogoliubov Fermi surfaces (TBFSs), and Fulde-Ferrell-Larkin-Ovchinnikov (FFLO) states with finite center-of-mass momentum. Among these, the FFLO states and TBFSs exemplify two unconventional forms of superconductivity. Considering the simplicity of this model, with only one band, zero net magnetization, and $s$-wave paring, the emergence of both unconventional phases underscores the pivotal role of altermagnetic spin splitting in enabling exotic pairing phenomena. This analytical study not only offers a valuable benchmark for future numerical simulations, but also provides a concrete experimental roadmap for realizing FFLO states and TBFSs in altermagnets.
I Introduction
The Fulde-Ferrell-Larkin-Ovchinnikov (FFLO) state (Fulde1964, ; Larkin1964, ) was first proposed in the 1960s in the context of Bardeen-Schrieffer-Cooper (BCS) superconductivity (BCS1957, ) under strong spin-exchange fields. In the presence of spin-population imbalance, Cooper pairs acquire a finite center-of-mass momentum $\bf q$} to minimize the free energy, therefore the translational symmetry of the homogeneous superfluidity (SF) is broken. Over the past six decades, extensive efforts have been devoted to realizing the FFLO phase across a wide range of platforms, including high-density quark matter (Casalbuoni2004, ; Alford2008, ), ultracold Fermi gases Hu2006 ; Hu2007 ; Liu2007 ; Radzihovsky2009 ; Liao2010 ; Liu2013 ; Cao2014 ; Takahashi2014 ; Sheehy2015 ; Wang2018 ; Kawamura2022 , and electronic materials (Shimahara1994, ; Pickett1999, ; Radovan2003, ; Uji2006, ; Matsuda2007, ; Kenzelmann2008, ; Mayaffre2014, ; Cho2017, ; Liu2017, ; Zhao2023, ; Wan2023, ). However, conclusive evidence-namely, spatial modulation of either the phase or amplitude of the SF order parameter-remains elusive.
Altermagnetism (AM) (Smejkal2022-1, ), a novel type of collinear magnetism beyond the traditional ferromagnetism (FM) and antiferromagnetism (AFM), has recently attracted considerable attention (Smejkal2022-2, ; Mazin2022, ; Bai2024, ; Song2025, ; Guo2025, ; Jungwirth2025-1, ; Jungwirth2025-2, ). As a hybrid of FM and AFM, AM features compensated collinear spin configurations in real space (akin to conventional AFM) while exhibits non-relativistic spin splitting (NRSS) in momentum space (reminiscent of FM) (Lee2024, ; Souma2024, ; Reimers2024, ; Ding2024, ; Krempasky2024, ; Yang2025, ; Leiviska2024, ; Han2024, ; Zhou2025, ; Jiang2025, ; Zhang2025, ). Combing both merit of FM and AFM, AM offers exciting opportunities for novel applications in spintronics (Naka2019, ; Smejkal2020, ; Reichlova2024, ), quantum geometry (Fang2024, ; Gong2024, ; LiY2024-1, ; Hernandez2025, ; Liu2025, ), and superconductivity (Mazin2022-2, ; Zhu2023, ; Brekke2023, ; Bose2024, ; Wu2025, ; LiY2024-2, ; deCarvalho2024, ; Chakraborty2024-2, ; LiY2023, ; Ghorashi2024, ; Hodge2025, ; Ouassou2023, ; Lu2024, ; Fukaya2025, ; Banerjee2024, ; Chakraborty2024-3, ; Papaj2023, ; Sun2023, ; Das2024, ; Wei2024, ; Carvalho2024, ; Parthenios2025, ).
AM also provides a promising platform for the FFLO state for two main reasons. First, FFLO phases typically require a nonzero spin imbalance, which is usually achieved via external Zeeman fields. However, in conventional superconductors, the orbital effect induced by magnetic fields can suppress the superconductivity before the FFLO emerges (Maki1966, ; Gruenberg1966, ). In contrast, AM naturally hosts NRSS without the need of external magnetic field, thereby avoiding detrimental orbital effects. Second, conventional FFLO state involves spontaneously breaking of the continuous $U(1)$ rotational symmetry of the Fermi surfaces, which, according to Nambu-Goldstone theorem (Nambu1960, ; Goldstone1961, ; Goldstone1962, ), leads to gapless modes associated with fluctuations between equivalent $\bf q$}. In AM, the continuous rotational symmetry is reduced to a discrete one Wu2007 , significantly suppressing such fluctuations. These unique advantages have motivated several theoretical investigations of FFLO phases in altermagnetic metals (Feiguin2009, ; Soto-Garrido2014, ; Gukelberger2014, ; Sumita2023, ; Chakraborty2024-1, ; ZhangS2024, ; Hong2025, ; Hu2025-1, ; Hu2025-2, ).
Despite these efforts, the existence of FFLO state in two-dimensional (2D) $d$-wave AM Fermi gas with $s$-wave pairing remains debated. Using a continuum model, Soto-Garrido and co-workers (Soto-Garrido2014, ) reported the absence of FFLO state in a 2D $d_{x^{2}-y^{2}}$-wave AM Fermi gas. This conclusion is corroborated by Feiguin et al. (Feiguin2009, ) and Chakraborty et al. (Chakraborty2024-1, ) through a square lattice model. Nevertheless, more recent studies challenge this result. Investigating a similar lattice model to that of Chakraborty et al. (Chakraborty2024-1, ), Hong and co-workers (Hong2025, ) found that FFLO state does emerge under purely $s$-wave pairing. Subsequently, Hu and co-workers (Hu2025-1, ) conducted a numerical calculation on the continuum model close to Soto-Garrido et al. (Soto-Garrido2014, ) and confirmed the existence of FFLO state, aligning with the findings of Hong et al. (Hong2025, ). These conflicting reports raise a fundamental question: can 2D $d$-wave AM Fermi gas supports the FFLO state driven solely by $s$-wave pairing?
In this work, we give a definitive and affirmative answer by analytically determining the ground-state phase diagram. We consider two physically relevant situations: (i) fixed $\mu/\varepsilon_{B}$, where $\mu$ is the total chemical potential and $\varepsilon_{B}$ is the two-body binding energy which is experimentally controllable in the ultracold atom systems; and (ii) fixed total particle number $N$, which is appropriate for describing electronic superconductors. In both situations, we demonstrate that FFLO phases do indeed exist. What’s more, another exotic unconventional form of superconductivity - nodal SF hosting topological Bogoliubov Fermi surface (TBFS) - also emerges in both situations.
The rest of the paper is organized as follows. In the next section (Sec. II), we outline the model Hamiltonian that describes a $d$-wave altermagnetic metal. In Sec. III, we briefly discuss how to solve the model Hamiltonian by using the standard Bogoliubov de-Gennes (BdG) mean-field theory, with an inhomogeneous FF order parameter. In Sec. IV, we present the phase diagram between normal (N) phase, nodal SF, and SF. The nontrivial $\mathbb{Z}_{2}$ topology of BFS is characterized by the Pfaffian of the BdG Hamiltonian. In Sec. V, we discuss in details the FFLO phase boundary driven by the NRSS. Finally, we conclude in Sec. VI and present some outlooks for future works.
To simplify the notation, the Fermi wavevector $k_{F}$ and Fermi energy $\varepsilon_{F}$ are taken as the units for wavevector and energy, respectively. This means that we can directly set $2m=\hbar=k_{B}=k_{F}=\varepsilon_{F}=1$.
II Model Hamiltonian
The model Hamiltonian describing two-dimensional spin-1/2 Fermion gas is:
${\mathscr{H}}$ $=$ $\int d{\bf x}({\cal H}_{0}+{\cal H}_{\textrm{int}}),$ (1)
${\cal H}_{0}$ $=$ $\sum_{\sigma=\uparrow,\downarrow}\psi_{\sigma}^{\dagger}\left({\bf x}\right)\left[\hat{\xi}_{{\bf k}}+s(\sigma)\hat{J}_{\mathbf{k}}\right]\psi_{\sigma}\left({\bf x}\right),$ (2)
${\cal H}_{\textrm{int}}$ $=$ $g\psi_{\uparrow}^{\dagger}\left({\bf x}\right)\psi_{\downarrow}^{\dagger}\left({\bf x}\right)\psi_{\downarrow}\left({\bf x}\right)\psi_{\uparrow}\left({\bf x}\right).$ (3)
For the non-interacting part ${\cal H}_{0}$, $\hat{\varepsilon}_{{\bf k}}\equiv-{\bf\nabla}^{2}$ is the operator for kinetic energy where $\hat{k}_{x}=-i\partial_{x}$, $\hat{k}_{y}=-i\partial_{y}$. $\hat{\xi}_{{\bf k}}\equiv\hat{\varepsilon}_{{\bf k}}-\mu$ is the modified kinetic energy operator where $\mu$ is the chemical potential. The chemical potential is tuned to yield the given total particle density $n=n_{\uparrow}+n_{\downarrow}$. $\hat{J}_{\mathbf{k}}\equiv t_{\textrm{AM}}2\hat{k}_{x}\hat{k}_{y}$ where $s(\uparrow)=+1$ and $s(\downarrow)=-1$ is the $d_{xy}$-wave altermagnetic spin splitting, whose strength is controlled by the dimensionless parameter $t_{\textrm{AM}}$ (without the loss of generality, $t_{\textrm{AM}}$ is set to be positive here). In 2D, there are two independent $d$-wave spin splitting: $d_{xy}$ and $d_{x^{2}-y^{2}}$. As $d_{x^{2}-y^{2}}$ can be obtained from $d_{xy}$ by $\pi/4$ rotation, the results obtained here can be generalized to $d_{xy}$-wave altermagnetic spin splitting straightforwardly. Without $\hat{J}_{\mathbf{k}}$, the spin degenerated Fermi surface is a circle with radius $\sqrt{\mu}$, which conserves $U(1)$ rotational symmetry (see the dashed circle in Fig. 1(a)). The introduction of $\hat{J}_{\mathbf{k}}$ distorts the Fermi surface into two orthogonal ellipses (denoted as Fermi ellipse here) with equal semi-major axis $a\equiv\sqrt{\mu/(1-t_{\textrm{AM}})}$ and semi-minor axis $b\equiv\sqrt{\mu/(1+t_{\textrm{AM}})}$ (see Fig. 1(a)), in this regard, $t_{\textrm{AM}}$ also measures the eccentricity of the Fermi ellipses at a given $\mu$. These two Fermi ellipses together exhibit $d_{xy}$-wave form (see Fig. 1(a)): spin up and down ellipse has major axis along $\left(1,-1\right)$ and $\left(1,1\right)$ direction, respectively. Since the two Fermi ellipses always have the same area even with finite momentum boost in the FFLO phase, the system remains spin-population balanced in the whole phase diagram. Such a circle-to-ellipse distortion breaks the continuous $U(1)$ to the discrete $C_{4z}$ symmetry, leaving four degeneracies along both $(1,0)$ and $(0,1)$ directions. As mentioned before, such continuous symmetry breaking would stabilize the FFLO state.
For the inter-particle interaction ${\cal H}_{\textrm{int}}$, we consider a short-range $s$-wave contact potential. Such a consideration is straightforward as pair states like $\left|\mathbf{k},\uparrow\right\rangle$ and $\left|-\mathbf{k},\downarrow\right\rangle$ in Fig. 1(a) are susceptible to $s$-wave pairing according to BCS theory (BCS1957, ). In two dimensions, its interaction strength $g$ can be conveniently regularized by using a two-body binding energy $\varepsilon_{B}$ (He2015, ):
$\frac{1}{g}=-\frac{1}{\mathcal{A}}\sum_{\mathbf{k}}\frac{1}{2\varepsilon_{{\bf k}}+\varepsilon_{B}},$ (4)
where $\mathcal{A}$ is the area of the 2D Fermi system.
III Mean-field theory
To search for the FF superconductivity, we assume a FF-like order parameter $\Delta({\bf x})=-g\left\langle\psi_{\downarrow}\left({\bf x}\right)\psi_{\uparrow}\left({\bf x}\right)\right\rangle=\Delta\exp[i\mathbf{q}\cdot\mathbf{x}]$ at positive $\mathbf{x}$ with the center-of-mass momentum $\mathbf{q}$, so $\left|\Delta({\bf x})\right|^{2}=\Delta^{2}$ (below we take $\Delta$ positive) and consider the standard mean-field decoupling of the interaction Hamiltonian density (Hu2006, ; Liu2013, ),
${\cal H}_{\textrm{int}}\simeq-\left[\Delta({\bf x})\psi_{\uparrow}^{\dagger}\left({\bf x}\right)\psi_{\downarrow}^{\dagger}\left({\bf x}\right)+\text{H.c.}\right]-\frac{1}{g}\Delta^{2}.$ (5)
In the framework of BdG mean-field theory, the total Hamiltonian can be rewritten into the form (Hu2006, ; Liu2013, ),
${\cal\mathscr{H}}_{\textrm{MF}}=\int d{\bf x}\Phi^{\dagger}\left({\bf x}\right){\cal H}_{\textrm{BdG}}\Phi\left({\bf x}\right)-\frac{\Delta^{2}}{g}\mathcal{A}+\mathcal{E}_{0}^{\prime},$ (6)
where $\Phi\left({\bf x}\right)\equiv[\psi_{\uparrow}\left({\bf x}\right),\psi_{\downarrow}^{\dagger}\left({\bf x}\right)]^{T}$ is the Nambu spinor and
${\cal H}_{\textrm{BdG}}\equiv\left[\begin{array}[]{cc}\hat{\xi}_{{\bf k}}+\hat{J}_{\mathbf{k}}&-\Delta\left({\bf x}\right)\\ -\Delta^{*}\left({\bf x}\right)&-\left(\hat{\xi}_{{\bf k}}-\hat{J}_{\mathbf{k}}\right)\end{array}\right],$ (7)
and $\mathcal{E}_{0}^{\prime}=\sum_{\mathbf{k}}(\hat{\xi}_{{\bf k}}-\hat{J}_{\mathbf{k}})$ is an energy shift (to be specified below), due to the abnormal order in the field operators $\psi_{\downarrow}(\mathbf{x})$ and $\psi_{\downarrow}^{\dagger}({\bf x})$ taken in the Nambu spinor representation.
Figure 1: (a) $d_{xy}$-wave spin-splitting by plotting $\xi_{{\bf k}}+s(\sigma)J_{{\bf k}}=0$, here $\mu=1$. $t_{\textrm{AM}}=0$ gives the dashed circle, and $t_{\textrm{AM}}=0.5$ gives the red and blue ellipses, corresponding to spin up and down channel, respectively. (b) Contour plot of $E_{{\bf k\pm}}=0$, here $\mu=1$, $t_{\textrm{AM}}=0.5$, and $\Delta=0$. The dash line corresponds to $\sum_{\eta}E_{{\bf k\eta}}=0$. Cyan and blue colored regions correspond to negative $E_{{\bf k+}}$ and $E_{{\bf k-}}$. (c) Curve for $E_{k}$ at different $t_{\textrm{AM}}$ values, here $\mu=1,\Delta=0.2$. (d) Plot of $\bar{t}_{\textrm{AM}}$ (see Eq. (42) for definition) as a function of $\Delta/\mu$. (e)-(f) Anisotropic Bogoliubov quasiparticle energy $E_{+}(\mathbf{k})$ and $E_{-}(\mathbf{k})$. The while solid line represent the Bogoliubov Fermi surface. Here $\mu=1$, $\Delta=0.2$, and $t_{\textrm{AM}}=0.5$, which corresponds to the red dot in (d).
III.1 Bogoliubov quasiparticles
As the FF order parameter takes a plane-wave form, it is convenient to diagonalize the mean-field Hamiltonian in momentum space, with the following Bogoliubov equations (Liu2013, ),
${\cal H}_{\textrm{BdG}}\Phi_{{\bf k}}\left({\bf x}\right)=E_{{\bf k}}\Phi_{{\bf k}}\left({\bf x}\right),$ (8)
where
$\Phi_{{\bf k}}\left({\bf x}\right)\equiv\left[\begin{array}[]{c}u_{{\bf k}\uparrow}e^{+i{\bf q}{\bf x}/2}\\ v_{{\bf k}\downarrow}e^{-i{\bf q}{\bf x}/2}\end{array}\right]e^{i{\bf kx}}$ (9)
and $E_{{\bf k}}$ are the wavefunction and energy of the Bogoliubov quasiparticles, respectively. With this wavefunction, the energy shift is $\mathcal{E}_{0}^{\prime}=\sum_{\mathbf{k}}(\xi_{{\bf k}-\mathbf{q}/2}-J_{\mathbf{k}-\mathbf{q}/2})$, where $\xi_{\mathbf{k}}\equiv k^{2}-\mu$, and $J_{\mathbf{k}}\equiv t_{\textrm{AM}}2k_{x}k_{y}$. Then the Bogoliubov equations becomes,
$\left[{\cal H}_{\textrm{BdG}}\right]\left[\begin{array}[]{c}u_{{\bf k}\uparrow}\\ v_{{\bf k}\downarrow}\end{array}\right]=E_{{\bf k}}\left[\begin{array}[]{c}u_{{\bf k}\uparrow}\\ v_{{\bf k}\downarrow}\end{array}\right],$ (10)
where $\left[{\cal H}_{\textrm{BdG}}\right]$ is a 2 by 2 matrix given by,
$\left[\begin{array}[]{cc}\xi_{{\bf k}+\mathbf{q}/2}+J_{\mathbf{k}+\mathbf{q}/2}&-\Delta\\ -\Delta&-(\xi_{{\bf k}-\mathbf{q}/2}-J_{\mathbf{k}-\mathbf{q}/2})\end{array}\right].$ (11)
$\left[{\cal H}_{\textrm{BdG}}\right]$ can be diagonalized with eigenvectors $u_{{\bf k}\uparrow}^{(\eta)},v_{{\bf k}\downarrow}^{(\eta)}]^{T}$ and corresponding eigenvalues $\eta E_{{\bf k}\eta}$, labeled by the branch index $\eta=\pm$. Explicitly, we have the Bogoliubov quasiparticle energy spectrum,
$E_{\mathbf{k}\pm}=\sqrt{A_{\mathbf{k}}^{2}+\Delta^{2}}\pm B_{\mathbf{k}},$ (12)
where
$A_{\mathbf{k}}$ $\equiv$ $\xi_{{\bf k}}+\frac{{\bf q}^{2}}{4}+t_{\textrm{AM}}(k_{x}q_{y}+k_{y}q_{x}),$ (13)
$B_{\mathbf{k}}$ $\equiv$ ${\bf k}\cdot{\bf q}+J_{\mathbf{k}}+\frac{t_{\textrm{AM}}}{2}q_{x}q_{y}.$ (14)
Here, as usual we have reversed the sign for the eigenvalue $E_{\mathbf{k}-}$, in order to write the corresponding field operators of Bogoliubov quasiparticles $\alpha_{\mathbf{k}-}$ and $\alpha_{\mathbf{k}-}^{\dagger}$ in the normal order. This introduces another energy shift $\mathcal{E}_{0}^{\prime\prime}=-\sum_{\mathbf{k}}E_{\mathbf{k}-}$ in the mean-field Hamiltonian. When $t_{\textrm{AM}}=0$, our quasiparticle energy spectrum reduces to the results of Ref. (Sheehy2015, ). In the end, we obtain the diagonalized mean-field Hamiltonian,
${\cal\mathscr{H}}_{\textrm{MF}}=\sum_{{\bf k},\eta=\pm}E_{{\bf k}\eta}\alpha_{{\bf k\eta}}^{\dagger}\alpha_{{\bf k}\eta}-\frac{\Delta^{2}}{g}\mathcal{A}+\mathcal{E}_{0},$ (15)
where
$\mathcal{E}_{0}=\mathcal{E}_{0}^{\prime}+\mathcal{E}_{0}^{\prime\prime}=\sum_{\mathbf{k}}\left(A_{\mathbf{k}}-\sqrt{A_{\mathbf{k}}^{2}+\Delta^{2}}\right).$ (16)
III.2 Thermodynamic potential and ground state energy
From the diagonalized Hamiltonian, we can straightforwardly write down the grand thermodynamic potential,
$\Omega=-\frac{\Delta^{2}}{g}\mathcal{A}+\mathcal{E}_{0}-T\sum_{{\bf k},\eta=\pm}\ln\left(1+e^{-\frac{E_{{\bf k\eta}}}{T}}\right),$ (17)
where $T$ is the temperature, and the last term is the standard contribution of non-interacting fermionic particles to the thermodynamic potential.
In the ground state at zero temperature, referred to as the ground-state energy, $\Omega$ takes the following form,
$E_{G}=-\frac{\Delta^{2}}{g}\mathcal{A}+\mathcal{E}_{0}+\sum_{{\bf k},\eta=\pm}E_{{\bf k\eta}}\Theta\left(-E_{{\bf k\eta}}\right),$ (18)
where $\Theta(x)$ is the Heaviside step function. Converting the momentum sum to an integration, the ground-state energy density becomes:
$\begin{split}\frac{E_{G}}{\mathcal{A}}&=\int\frac{d{\bf k}}{\left(2\pi\right)^{2}}\left[\sum_{\eta=\pm}E_{{\bf k\eta}}\Theta\left(-E_{{\bf k\eta}}\right)\right.\\ &\left.+\frac{\Delta^{2}}{2\varepsilon_{{\bf k}}+\varepsilon_{B}}+A_{\mathbf{k}}-\sqrt{A_{\mathbf{k}}^{2}+\Delta^{2}}\right],\end{split}$ (19)
where the renormalization condition Eq. (4) is used. At a given chemical potential, the phase boundary between SF and N phase is obtained by equating the ground-state energies of these two phases:
$E_{G,\textrm{SF}}=E_{G,\textrm{N}}.$ (20)
III.3 Gap equation
During the phase transition from N to FFLO SF phase, $E_{G}$ should decrease as the onset of $\Delta$. The minimization of $E_{G}$ with respect to $\Delta$ defines a function $S\left(t_{\textrm{AM}},{\bf q},\Delta\right)$:
$2\Delta S\left(t_{\textrm{AM}},{\bf q},\Delta\right)=\frac{\partial E_{G}}{\partial\Delta},$ (21)
with the explicit expression
$\frac{S\left(t_{\textrm{AM}},{\bf q},\Delta\right)}{\mathcal{A}}=\int\frac{d{\bf k}}{\left(2\pi\right)^{2}}\left[\frac{1-\sum_{\eta}\Theta\left(-E_{{\bf k\eta}}\right)}{\sum_{\eta}E_{{\bf k\eta}}}-\frac{1}{2\varepsilon_{{\bf k}}+\varepsilon_{B}}\right].$ (22)
Equilibrium FFLO states satisfy $S\left(t_{\textrm{AM}},{\bf q},\Delta\right)=0$, which is also known as the gap equation. In the following, this gap equation is utilized to find the phase boundary between N and FFLO SF phase. We note that, $S(t_{\textrm{AM}},{\bf q},\Delta)$ can alternatively be understood as the inverse vertex function (i.e., the Cooper-pair propagator in the normal state with frequency $\omega=\Delta$) in the ladder approximation Sheehy2015 ; Hu2025-2 . Therefore, the condition $S(t_{\textrm{AM}},{\bf q},0)=0$ is precisely the well-known Thouless criterion for the onset of the N to SF phase transition.
IV Phase diagram at ${\bf q}={\bf 0}$
In this section, we determine the phase boundaries assuming ${\bf q}={\bf 0}$ pairing, which amounts to neglecting the possibility of the FFLO state. Within this approximation, $A_{{\bf k}}=\xi_{{\bf k}}$ and $E_{{\bf k\pm}}=\sqrt{\xi_{{\bf k}}^{2}+\Delta^{2}}\pm J_{\mathbf{k}}$. In our terminology, $E_{{\bf k\pm}}$ is non-negative. If $E_{{\bf k\pm}}>0$ for all $\bf k$}, such a SF phase is called BCS SF. If $E_{{\bf k\pm}}=0$ at certain $\bf k$}, then the Bogoliubov quasiparticle has a “Fermi surface” structure, which is also known as BFS (Volovik1993, ). As discussed below, we have nodal SF with BFS protected by $\mathbb{Z}_{2}$ topological invariant. In the following, we first investigate the BCS SF phase.
IV.1 BCS SF phase
Since $E_{{\bf k\pm}}$ is positive, the Heaviside step function in Eq. (22) vanishes and $S\left(t_{\textrm{AM}},{\bf 0},\Delta\right)/\mathcal{A}$ becomes
$\begin{split}\frac{S\left(t_{\textrm{AM}},{\bf 0},\Delta\right)}{\mathcal{A}}&=\int\frac{d{\bf k}}{\left(2\pi\right)^{2}}\left[\frac{1}{2\sqrt{\xi_{{\bf k}}^{2}+\Delta^{2}}}-\frac{1}{2\varepsilon_{{\bf k}}+\varepsilon_{B}}\right],\\ &=\frac{1}{8\pi}\ln\left[\frac{\varepsilon_{B}}{\sqrt{\mu^{2}+\Delta^{2}}-\mu}\right].\end{split}$ (23)
And the gap equation leads to the stationary pairing amplitude
$\Delta=\sqrt{\varepsilon_{B}\left(\varepsilon_{B}+2\mu\right)},$ (24)
so we must have $\mu>-\varepsilon_{B}/2$ for a stable SF phase.
The total particle number of the SF state at a fixed $\mu$ is given by calculating $N=-\frac{E_{G,\textrm{SF}}}{\partial\mu}$, yielding, for the total density
$n=\frac{1}{2\pi}\left(\mu+\sqrt{\mu^{2}+\Delta^{2}}\right).$ (25)
Inserting Eq. (24) into Eq. (25) gives results for $\Delta$ and $\mu$ for a system at a fixed total particle number
$\begin{split}\frac{\Delta}{\varepsilon_{F}}&=\sqrt{\frac{\varepsilon_{B}}{\varepsilon_{F}}},\\ \frac{\mu}{\varepsilon_{F}}&=1-\frac{1}{2}\frac{\varepsilon_{B}}{\varepsilon_{F}}.\end{split}$ (26)
Here, we want to emphasize that $t_{\textrm{AM}}$ does not appear in Eq. (24), Eq. (25), and Eq. (26). This is in agreement with the earlier result in Ref. (Sheehy2015, ), where a small magnetic field does not appear in the stationary pairing amplitude of a BCS SF.
IV.2 Fixed chemical potential
At the SF phase, the effect of NRSS disappears. The integration over momentum and the insertion of the stationary gap-equation solution leads to
$\frac{E_{G,\textrm{SF}}}{\mathcal{A}}=-\frac{1}{16\pi}(\varepsilon_{B}+2\mu)^{2}$ (27)
The effect of NRSS does appear in the normal phase. By setting $\Delta=0$ in Eq. (19), we find that,
$\frac{E_{G,\textrm{N}}}{\mathcal{A}}=\int\frac{d{\bf k}}{\left(2\pi\right)^{2}}\left[\sum_{\eta=\pm}E_{{\bf k\eta}}\Theta\left(-E_{{\bf k\eta}}\right)+\left(\xi_{{\bf k}}-\left|\xi_{{\bf k}}\right|\right)\right],$ (28)
where $E_{{\bf k\eta}}=\left|\xi_{{\bf k}}\right|\pm J_{{\bf k}}$, i.e. the energies of Bogoliubov quasiparticles reduce to the standard single-particle excitation energies.
The integration over the second term in the above equation is straightforward:
$\int\frac{d{\bf k}}{\left(2\pi\right)^{2}}\left(\xi_{{\bf k}}-\left|\xi_{{\bf k}}\right|\right)=-\frac{1}{4\pi}\mu^{2}$ (29)
Next we integrate over the first term. Introducing polar coordinate ${\bf k}=[k\cos\left(\phi\right),k\sin\left(\phi\right)]$ and setting $E_{{\bf k\eta}}=0$, the boundary $k_{\pm}(\phi)$ is already shown in Fig. 1(a) with the following expression:
$k_{\pm}(\phi)=\sqrt{\frac{\mu}{1\mp t_{\textrm{AM}}\sin\left(2\phi\right)}}.$ (30)
Here we see $t_{\textrm{AM}}<1$. $E_{{\bf k+}}$ ($E_{{\bf k-}}$) becomes negative only when $\phi\in\left[\frac{\pi}{2},\pi\right]\cup\left[\frac{3\pi}{2},2\pi\right]$ ($\left[0,\frac{\pi}{2}\right]\cup\left[\pi,\frac{3\pi}{2}\right]$), which is marked as the cyan (blue) region in Fig. 1(b). Due to the $C_{4z}$ symmetry, it is clear that $\int\frac{d{\bf k}}{\left(2\pi\right)^{2}}E_{{\bf k+}}\Theta\left(-E_{{\bf k+}}\right)$ and $I_{1}\equiv\int\frac{d{\bf k}}{\left(2\pi\right)^{2}}E_{{\bf k-}}\Theta\left(-E_{{\bf k-}}\right)$ are the same, so only $I_{1}$ is evaluated here. What is more, the contribution from $\phi\in\left[0,\frac{\pi}{2}\right]$ and $\phi\in\left[\pi,\frac{3\pi}{2}\right]$ are also the same, therefore, we have
$I_{1}=\frac{2}{\left(2\pi\right)^{2}}\int_{0}^{\pi/2}d\phi\int_{k_{-}(\phi)}^{k_{+}(\phi)}\left(\left|\xi_{{\bf k}}\right|-J_{{\bf k}}\right)kdk.$ (31)
By noticing that $\left|\xi_{{\bf k}}\right|=-\xi_{{\bf k}}$ for $k\in\left[k_{-}(\phi),\sqrt{\mu}\right]$ and $\xi_{{\bf k}}$ for $k\in\left[\sqrt{\mu},k_{+}(\phi)\right]$, the integration over $k$ can be calculated as
$2\int_{k_{-}(\phi)}^{k_{+}(\phi)}\left(\left|\xi_{{\bf k}}\right|-J_{{\bf k}}\right)kdk=-\frac{\left[t_{\textrm{AM}}\sin(2\phi)\right]^{2}}{1-\left[t_{\textrm{AM}}\sin(2\phi)\right]^{2}}\mu^{2},$ (32)
and $I_{1}$ becomes the integration over $\phi$ only:
$\begin{split}I_{1}&=-\frac{\mu^{2}}{\left(2\pi\right)^{2}}\int_{0}^{\pi/2}d\phi\left[\frac{1}{1-\left[t_{\textrm{AM}}\sin(2\phi)\right]^{2}}-1\right],\\ &=-\frac{\mu^{2}}{\left(2\pi\right)^{2}2t_{\textrm{AM}}^{2}}\int_{0}^{2\pi}\frac{1}{\frac{2-t_{\textrm{AM}}^{2}}{t_{\textrm{AM}}^{2}}+\cos(\phi)}d\phi+\frac{\mu^{2}}{8\pi}.\end{split}$ (33)
With $t_{\textrm{AM}}<1$, we have $\frac{2-t_{\textrm{AM}}^{2}}{t_{\textrm{AM}}^{2}}>1$, the standard integration gives us
$\int_{0}^{2\pi}\frac{1}{\frac{2-t_{\textrm{AM}}^{2}}{t_{\textrm{AM}}^{2}}+\cos(\phi)}d\phi=\frac{2\pi}{\sqrt{\left(\frac{2-t_{\textrm{AM}}^{2}}{t_{\textrm{AM}}^{2}}\right)^{2}-1}},$ (34)
and finally $I_{1}$ is expressed as:
$I_{1}=\frac{\mu^{2}}{8\pi}\left(1-\frac{1}{\sqrt{1-t_{\textrm{AM}}^{2}}}\right).$ (35)
Therefore, we obtain
$\frac{E_{G,\textrm{N}}}{\mathcal{A}}=-\frac{1}{4\pi}\frac{\mu^{2}}{\sqrt{1-t_{\textrm{AM}}^{2}}}.$ (36)
It is worth noting that the area of the ellipse in Fig. 1(a) is given by $\pi ab=\pi\mu/\sqrt{1-t_{\textrm{AM}}^{2}}$, therefore, the ground-state energy density of the N phase is proportional to the area of the ellipse. Such a geometric meaning can also be understood by setting $t_{\textrm{AM}}=0$, the above ground-state energy density becomes the well-known result $E_{G,\textrm{N}}/\mathcal{A}=-\mu^{2}/(4\pi)$.
By equating $E_{G,\textrm{SF}}$ and $E_{G,\textrm{N}}$, we have
$(\varepsilon_{B}+2\mu)^{2}=\frac{\left(2\mu\right)^{2}}{\sqrt{1-t_{\textrm{AM}}^{2}}},$ (37)
and the critical $t_{\textrm{AM}}$ separating SF and N phase is given by:
$t_{\textrm{AM, c}}=\sqrt{1-\left(\frac{2\mu/\varepsilon_{B}}{1+2\mu/\varepsilon_{B}}\right)^{4}}.$ (38)
This determines the black solid curve in Fig. 2(a). When $t_{\textrm{AM}}$ is small, i.e. anisotropic spin splitting is weak, we have SF phase, even to the high chemical potential regime. Such a feature is in accordance with the earlier result that SF phase exist at high chemical potential regime when Zeeman field is small Sheehy2015 . Such SF becomes unstable when the anisotropic NRSS becomes large.
Figure 2: (a) Phase diagram at fixed chemical potential $\mu$ (normalized to the two-body binding energy $\varepsilon_{B}$) versus NRSS $t_{\textrm{AM}}$. (b) Phase diagram at fixed total particle number $N$, as a function of the binding energy $\varepsilon_{B}$ normalized to the Fermi energy $\varepsilon_{F}$. The black curves in both figures show the Pauli-limit, upper altermagnetic coupling strength, where a SF with zero center-of-mass momentum $\mathbf{q}=0$ becomes unstable towards a normal state, when the possibility of an inhomogeneous FFLO is not considered. In the weak-coupling regime (i.e., for large $\mu/\varepsilon_{B}$ or small $\varepsilon_{B}/\varepsilon_{F}$), those two black curves provide an excellent estimate for the phase boundary between a TBFS phase and a FFLO state, within the superfluid domain.
IV.3 Fixed total particle number
After determining the phase boundary at fixed chemical potential, our next task is to consider the case of a fixed total particle number. By approaching the phase transition from the N phase, $\mu=\sqrt{1-t_{\textrm{AM}}^{2}}\varepsilon_{F}$ according to Eq. (36). Then Eq. (38) leads to
$t_{\textrm{AM, c}}=\sqrt{\frac{1}{2}+\frac{\varepsilon_{B}}{2\varepsilon_{F}}-\sqrt{\frac{1}{4}-\frac{\varepsilon_{B}}{2\varepsilon_{F}}}},$ (39)
which gives the black solid curve in Fig. 2(b) for $\varepsilon_{B}/\varepsilon_{F}<1/2$. When $t_{\textrm{AM}}=0$, arbitrarily weak attraction interactions are enough to drive the system into SF phase, as expected. The introduction of $t_{\textrm{AM}}$ will destabilize SF and the system transits to a normal state.
IV.4 Nodal SF
Now we study the nodal SF. Since the Bogoliubov quasiparticle energy is
$E_{k,\pm}\left(\phi\right)=\sqrt{\xi_{{\bf k}}^{2}+\Delta^{2}}\pm t_{\textrm{AM}}k^{2}\sin\left(2\phi\right),$ (40)
it is highly anisotropic: along $(1,0)$ and $(0,1)$ directions, i.e., $\phi=0$ and $\pi/2$, the pairing of the two states $\left|\mathbf{k},\uparrow\right\rangle$ and $\left|-\mathbf{k},\downarrow\right\rangle$ in Fig. 1(a) leads to an excitation energy $E_{{\bf k},\pm}=\sqrt{\xi_{{\bf k}}^{2}+\Delta^{2}}$ with minimum $\Delta$. Such a positive excitation energy follows the spirit of BCS theory BCS1957 . Along (1,1)(-1,1) and $(1,1)$ direction, i.e., $\phi=\frac{3\pi}{4}(\frac{7\pi}{4})$ and $\frac{\pi}{4}(\frac{5\pi}{4})$, $E_{\mathbf{k},+}$ and $E_{\mathbf{k},-}$ reach the same minimum value
$E_{\mathbf{k}}=\sqrt{\xi_{{\bf k}}^{2}+\Delta^{2}}-t_{\textrm{AM}}k^{2}.$ (41)
Fig. 1(c) shows the typical behavior of $E_{\mathbf{k}}$ with respect to different $t_{\textrm{AM}}$ values: $E_{\mathbf{k}}$ acquires a global minimum at $\xi_{{\bf k}}=\frac{t_{\textrm{AM}}}{\sqrt{1-t_{\textrm{AM}}^{2}}}\Delta$ with the value $E_{\mathbf{k},\textrm{min}}=\frac{1-t_{\textrm{AM}}}{\sqrt{1-t_{\textrm{AM}}^{2}}}\Delta-\mu t_{\textrm{AM}}$.
BFS exists when $E_{\mathbf{k},\textrm{min}}=0$ has real positive root ($\bar{t}_{\textrm{AM}}$) within $[0,1)$. As $E_{\mathbf{k},\textrm{min}}=0$ is equivalent to the cubic equation
$t_{\textrm{AM}}^{3}+t_{\textrm{AM}}^{2}+\left(\frac{\Delta}{\mu}\right)^{2}t_{\textrm{AM}}-\left(\frac{\Delta}{\mu}\right)^{2}=0,$ (42)
it is obvious that Eq. (42) has only one real root within the range $[0,1)$ as shown in Fig. 1(d). Therefore, when $t_{\textrm{AM}}>\bar{t}_{\textrm{AM}}$, $E_{k,\pm}\left(\phi\right)=0$ always have solutions and we have nodal SF with the typical BFS shown in Fig. 1(e)-(f). The closed BFS separates the whole momentum space into two regions, in the following the momentum inside (outside) BFS is marked as $\mathbf{k}_{-}$ ($\mathbf{k}_{+}$).
By inserting Eq. (24) into Eq. (42), we obtain the equation for determining $\bar{t}_{\textrm{AM}}$ at fixed $\mu$, which is displayed as the black dashed line in Fig. 2(a). At the same time, by plugging Eq. (26) in Eq. (42), we obtain the equation for determining $\bar{t}_{\textrm{AM}}$ at fixed total particle number, which is displayed as the black dashed line in Fig. 2(b). With $\bar{t}_{\textrm{AM}}<t_{\textrm{AM, c}}$, the nodal SF phase does exist in both fixed chemical potential and fixed total particle number cases. From Fig. 2(b), we know SF /nodal SF is preferred for a strong/weak pairing interactions, in accordance with Feiguin et al. (Feiguin2009, ).
IV.5 Topological BFS
The fact that the BFS shown in Fig. 1(e)-(f) shares the same dimension as the normal state Fermi surface indicates it has a topological origin. To reveal the topological invariant, here we hinges on the Pfaffian Pf($\mathbf{k}$) of Eq. 7. Since Eq. 7 preserves both charge conjuration and parity symmetry, as suggested by Agterberg and co-workers (Agterberg2017, ), we can always find a suitable basis to transform Eq. (7) into a screw form whose Pf($\mathbf{k}$) is given by
$\mathrm{Pf}(\mathbf{k})=\xi_{{\bf k}}^{2}+\Delta^{2}-\left[t_{\textrm{AM}}k^{2}\sin\left(2\phi\right)\right]^{2}.$ (43)
[…]
Here we would like to compare TBFS realized here with existing approaches. It is well known that BFS can not be realized in conventional 2D $s$-wave superconductors through external magnetic field alone, as the required magnetic field already drives the system into N phase before BFS can form. The replacement of the nodeless $s$-wave pairing by the nodal $d_{x^{2}-y^{2}}$-wave pairing has shown to enable TBFS (Yang1998, ; Setty2020-1, ). Alternatively, inducing supercurrent in $s$-wave superconductors with high Fermi velocity can also give rise to TBFS (Zhu2021, ). A further route involves multiband superconductors that preserve inversion symmetry but breaks time-reversal symmetry, as demonstrated in recent theoretical studies (Brydon2018, ; Lapp2020, ; Setty2020-2, ). In contrast to these approaches, AM introduces a new but simple mechanism for realizing TBFS. This scenario requires neither external magnetic field nor spin-orbit coupling, but only a single band and $s$-wave pairing, offering a minimal yet realistic platform for TBFS in superconducting systems.
V FFLO phase boundary driven by NRSS
Now we turn to the FFLO phase. To study the FFLO phase, we assume a continuous transition, with decreasing $t_{\textrm{AM}}$, from the N phase to the FFLO phase, occurring when the curvature of $E_{G}$ versus $\Delta$, at $\Delta=0$, becomes 0 for ${\bf q}\neq(0,0)$. This equals to $S(t_{\textrm{AM}},{\bf q},0)=0$ with
$\frac{S(t_{\textrm{AM}},{\bf q},0)}{\mathcal{A}}=\int\frac{d{\bf k}}{\left(2\pi\right)^{2}}\left[\frac{N({\bf k})}{2A_{{\bf k}}}-\frac{1}{2\varepsilon_{{\bf k}}+\varepsilon_{B}}\right]$ (44)
where we have defined the form factor $N({\bf k})=1-\Theta\left(-\left(A_{{\bf k}}+B_{{\bf k}}\right)\right)-\Theta\left(-\left(A_{{\bf k}}-B_{{\bf k}}\right)\right)$. It is clear that $N({\bf k})$ exhibits discontinuities whenever an argument of one of the Heaviside functions vanishes.
We write $S(t_{\textrm{AM}},{\bf q},0)/\mathcal{A}$ as the sum of the following three integrals,
$\begin{split}I_{2}&=\int\frac{d{\bf k}}{\left(2\pi\right)^{2}}\left[\frac{1}{2A_{{\bf k}}}-\frac{1}{2\varepsilon_{{\bf k}}+\varepsilon_{B}}\right],\\ I_{3}&=-\int\frac{d{\bf k}}{\left(2\pi\right)^{2}}\frac{\Theta\left[-\left(A_{{\bf k}}+B_{{\bf k}}\right)\right]}{2A_{{\bf k}}},\\ I_{4}&=-\int\frac{d{\bf k}}{\left(2\pi\right)^{2}}\frac{\Theta\left[-\left(A_{{\bf k}}-B_{{\bf k}}\right)\right]}{2A_{{\bf k}}}.\end{split}$ (45)
Among them, the evaluation of $I_{2}$ is straightforward:
$I_{2}=\frac{1}{8\pi}\left(\ln\frac{\varepsilon_{B}}{2}-\ln\left|\left(1-t_{\textrm{AM}}^{2}\right)\left(\frac{{\bf q}}{2}\right)^{2}-\mu\right|\right)$ (46)
The evaluation of $I_{3}$ and $I_{4}$ is highly nontrivial as the region of integration ($A_{{\bf k}}\pm B_{{\bf k}}<0$) is an ellipse while the integrand ($1/A_{{\bf k}}$) is circular. By coordinate transformation, $I_{3}$ and $I_{4}$ can be expressed as,
$\begin{split}I_{3}&=-\frac{ab}{2\left(2\pi\right)^{2}}\int_{0}^{2\pi}\int_{0}^{1}\frac{kdkd\phi}{f_{2}^{+}(\phi)k^{2}+f_{1}^{+}(\phi)k+f_{0}^{+}},\\ I_{4}&=-\frac{ab}{2\left(2\pi\right)^{2}}\int_{0}^{2\pi}\int_{0}^{1}\frac{kdkd\phi}{f_{2}^{-}(\phi)k^{2}+f_{1}^{-}(\phi)k+f_{0}^{-}},\end{split}$ (47)
where the coefficients are
$\begin{split}f_{2}^{\pm}(\phi)&=a^{2}\cos^{2}\left(\phi\right)+b^{2}\sin^{2}\left(\phi\right),\\ f_{1}^{+}(\phi)&=\frac{(1+t_{\textrm{AM}})(q_{y}-q_{x})}{\sqrt{2}}a\cos\left(\phi\right)\\ &+\frac{(-1+t_{\textrm{AM}})(q_{y}+q_{x})}{\sqrt{2}}b\sin\left(\phi\right),\\ f_{1}^{-}(\phi)&=\frac{(1+t_{\textrm{AM}})(q_{y}+q_{x})}{\sqrt{2}}a\cos\left(\phi\right)\\ &+\frac{(1-t_{\textrm{AM}})(q_{y}-q_{x})}{\sqrt{2}}b\sin\left(\phi\right),\\ f_{0}^{\pm}&=\frac{{\bf q}^{2}}{2}\pm t_{\textrm{AM}}q_{x}q_{y}-\mu.\end{split}$ (48)
There is a duality between $I_{3}$ and $I_{4}$. By changing $t_{\textrm{AM}}$ to $t_{\textrm{AM}}$ and $\phi$ to $\phi+\frac{3\pi}{2}$ simultaneously, then $a,b$ becomes $b,a$ and $\sin(\phi),\cos(\phi)$ becomes $\cos(\phi),\sin(\phi)$, therefore $f_{2}^{+}(\phi)$ becomes $f_{2}^{-}(\phi)$, $f_{1}^{+}(\phi)$ becomes $f_{1}^{-}(\phi)$, and $f_{0}^{+}$ becomes $f_{0}^{-}$. Since the translation of integral variable $\phi\rightarrow\phi+\frac{3\pi}{2}$ does not change the integration over $\phi$, $I_{3}$ and $I_{4}$ should be connected by $t_{\textrm{AM}}\rightarrow-t_{\textrm{AM}}$. Such a fact simplifies the integrations. In the following, we firstly evaluate $I_{4}$ by contour integration, then changing $t_{\textrm{AM}}$ to $t_{\textrm{AM}}$ to obtain $I_{3}$.
For $d_{xy}$-wave altermagnetic spin splitting, there are two characteristic high-symmetric directions for the FF momentum $\mathbf{q}$: along the $\left(1,0\right)$ direction (or $\left(0,1\right)$), and along the $\left(1,1\right)$ direction (or equivalently $\left(1,-1\right)$) according to the $C_{4z}$ rotational symmetry. In the following, we first analyze the case of the $\left(1,1\right)$ direction for concreteness, and then comment on the situation along the $\left(1,0\right)$ direction, which appears to be more difficult. As discussed in Appendix A, the angular dependence of the peak value of $S(t_{\textrm{AM}},{\bf q},0)/\mathcal{A}$ as a function of $q$ turns out to be weak. We thus anticipate that the analysis along the $\left(1,1\right)$ direction will provide a reasonably accurate phase boundary for the onset of the FFLO state.
Figure 3: (a)-(b) Sign distribution of $w_{+}^{2}-4$ and $w_{-}^{2}-4$. (c) When $w_{+}^{2}-4>0$ and $w_{-}^{2}-4>0$, $z_{1}$ and $z_{3}$ are within the unit circle while $z_{2}$ and $z_{4}$ are outside the unit circle. At $w_{+}^{2}-4=0$, $z_{1}$ and $z_{2}$ coincide with each other at $z=1$. After that, $w_{+}^{2}-4<0$ and $w_{-}^{2}-4>0$, we have (d) where $z_{1}$ and $z_{2}$ are on the unit circle, while $z_{3}$ ($z_{4}$) is inside (outside) the unit circle. When $w_{-}^{2}-4=0$, $z_{3}$ and $z_{4}$ coincide with each other at z=1z=-1. After that, $w_{+}^{2}-4<0$ and $w_{-}^{2}-4<0$, we have (e) where all four poles are on the unit circle. In (d)-(e), small arcs (blue and red solid line) are added to the unit circle to form a closed contour (solid line) to remove the poles on the unit circle.
V.1 Contour integration of $I_{4}$
There are two integral variables $k$ and $\phi$ in $I_{4}$. The integration order can be first $k$ then $\phi$, or first $\phi$ then $k$. We find that the second choice is much more convenient than the first one and thus adopt it. Since $\bf q$} is along the $(1,1)$ direction, we have $(q_{x},q_{y})=(\frac{q}{\sqrt{2}},\frac{q}{\sqrt{2}})$ ($q>0$ for simplicity) and the denominator of the integrand of $I_{4}$ is,
$\begin{split}D_{I_{4}}&=a^{2}k^{2}\cos^{2}(\phi)+b^{2}k^{2}\sin^{2}(\phi)+a(1+t_{\textrm{AM}})kq\cos(\phi)\\ &+\frac{1+t_{\textrm{AM}}}{2}q^{2}-\mu.\end{split}$ (49)
Applying Euler’s formula $\cos(\phi)\rightarrow\frac{1}{2}(z+\frac{1}{z})$, $D_{I_{4}}$ becomes
$D_{I_{4}}=\frac{(a^{2}-b^{2})k^{2}}{4}\left(z^{2}+\alpha z+\beta+\alpha z^{-1}+z^{-2}\right),$ (50)
where
$\begin{split}\alpha&=\frac{(1+t_{\textrm{AM}})^{2}\sqrt{1-t_{\textrm{AM}}}}{t_{\textrm{AM}}}\frac{Q}{k},\\ \beta&=\frac{2k^{2}+(1+t_{\textrm{AM}})^{2}(1-t_{\textrm{AM}})Q^{2}-2(1-t_{\textrm{AM}}^{2})}{t_{\textrm{AM}}k^{2}},\end{split}$ (51)
and $Q\equiv\frac{q}{\sqrt{\mu}}$ is the dimensionless momentum. The integral over $\phi$ now becomes a contour integral over the unit circle:
$\begin{split}I_{4}&=-\frac{1}{(2\pi)^{2}}\frac{2ab}{a^{2}-b^{2}}\int_{0}^{1}\frac{dk}{k}\oint_{|z|=1}f_{4}(z)dz,\\ f_{4}(z)&\equiv\frac{iz}{z^{4}+\alpha z^{3}+\beta z^{2}+\alpha z+1}.\end{split}$ (52)
The four poles (denoted as $z_{1/2/3/4}$) of $f_{4}(z)$ is determined by the equation $z^{2}+\alpha z+\beta+\alpha z^{-1}+z^{-2}=0$. Such an equation exhibits the reciprocal symmetry: if $z$ is a solution, so is $\frac{1}{z}$. Therefore, when considering relative position of $z_{1/2/3/4}$ with respect to the unit circle, there are three situations: 1) all $z_{1/2/3/4}$ are on the unit circle; 2) two poles are on the unit circle, one inside and one outside the unit circle. Here we always choose $z_{1}$ on the unit circle (so $z_{2}=\frac{1}{z_{1}}$ is also on the unit circle) and $z_{3}$ inside the unit circle (so $z_{4}=\frac{1}{z_{3}}$ is outside the unit circle). 3) two poles are inside the unit circle and two outside the unit circle. We choose $z_{1}$ and $z_{3}$ inside the unit circle, then $z_{2}=\frac{1}{z_{1}}$ and $z_{4}=\frac{1}{z_{3}}$ are outside the unit circle.
V.2 Positions of the four poles
Defining $w=z+\frac{1}{z}$, then $w^{2}+\alpha w+\beta-2=0$. The discriminant is $\mathcal{D}=(1-t_{\textrm{AM}})C(k)/(t_{\textrm{AM}}^{2}k^{2})$ with
$C(k)\equiv\left(1-t_{\textrm{AM}}^{2}\right)^{2}Q^{2}+8t_{\textrm{AM}}\left(1+t_{\textrm{AM}}-k^{2}\right).$ (53)
With $k\in[0,1]$ and $t_{\textrm{AM}}\in(0,1)$, we have $C(k)>0$ and $\mathcal{D}>0$. Then
$w_{\pm}=\frac{\sqrt{1-t_{\textrm{AM}}}}{2t_{\textrm{AM}}k}\left[-\left(1+t_{\textrm{AM}}\right)^{2}Q\pm\sqrt{C(k)}\right],$ (54)
so both $w_{+}$ and $w_{-}$ are purely real.
Suppose $z_{1/2}$ ($z_{3/4}$) corresponds to the solution $z=\frac{w_{+}\pm\sqrt{w_{+}^{2}-4}}{2}$ ($z=\frac{w_{-}\pm\sqrt{w_{-}^{2}-4}}{2}$). Since $w_{+}$ is purely real, if $w_{+}^{2}-4>0$, then we have $z_{1}=\frac{w_{+}-\sqrt{w_{+}^{2}-4}}{2}$ inside the unit circle and $z_{2}=\frac{w_{+}+\sqrt{w_{+}^{2}-4}}{2}$ outside the unit circle. Nevertheless, if $w_{+}^{2}-4<0$, then we have $z_{1/2}=\frac{w_{+}\pm i\sqrt{-w_{+}^{2}+4}}{2}$ both on the unit circle. The same analysis applies for $w_{-}$ and $z_{3/4}$.
The two discriminants are calculated as $w_{\pm}^{2}-4=(1+t_{\textrm{AM}})D_{\mp}(k)/(2t_{\textrm{AM}}^{2}k^{2})$ with
$\begin{split}D_{\pm}(k)&\equiv\left(1-t_{\textrm{AM}}^{4}\right)Q^{2}\pm\left(1-t_{\textrm{AM}}^{2}\right)Q\sqrt{C(k)}\\ &+4t_{\textrm{AM}}\left(1-t_{\textrm{AM}}-k^{2}\right).\end{split}$ (55)
If $k^{2}<1-t_{\textrm{AM}}$, then $D_{+}(k)>0$ and $w_{-}^{2}-4>0$. For the left cases, the condition $D_{\pm}(k)=0$ gives two critical $Q$ curves:
$\begin{split}Q_{c1}(k)&=\left|\sqrt{\frac{2-k^{2}}{1+t_{\textrm{AM}}}}-\sqrt{\frac{k^{2}}{1-t_{\textrm{AM}}}}\right|,\\ Q_{c2}(k)&=\sqrt{\frac{2-k^{2}}{1+t_{\textrm{AM}}}}+\sqrt{\frac{k^{2}}{1-t_{\textrm{AM}}}},\end{split}$ (56)
which are shown in Fig. 3(a)-(b). These two critical $Q$ values separate the $(k,Q)$-parameter space into different regions according to the sign of $w_{\pm}^{2}-4$. In Fig. 3(a) (Fig. 3(b)), the region marked by “+” sign means $w_{+}^{2}-4>0$ ($w_{-}^{2}-4>0$), while $w_{+}^{2}-4<0$ ($w_{-}^{2}-4<0$) in the region marked by “-” sign.
At given $(k,Q)$, there are three different possibilities according to sign of $w_{\pm}^{2}-4$: (a) both are positive, so poles $z_{1}$ and $z_{3}$ are inside the unit circle while $z_{2}$ and $z_{4}$ outside the unit circle as shown in Fig. 3(c); (b) $w_{+}^{2}-4$ is negative while $w_{-}^{2}-4$ is positive, so poles $z_{1}$ and $z_{2}$ are on the unit circle, while $z_{3}$ inside and $z_{4}$ outside the unit circle, as shown in Fig. 3(d); and (c) both are negative, so all four poles are on the unit circle, as shown in Fig. 3(e). Accordingly, the integration $I_{4}(k)\equiv\oint_{|z|=1}f_{4}(z)dz$ leads to three possibilities and in the following, we will use $I_{4}^{(i)}(k)$ ($i=1,2,3$) to represent these different situations.
V.3 Integrating over $\phi$
We first evaluate $I_{4}^{(1)}(k)$ where the poles are shown in Fig. 3(c). Since $z_{1}$ and $z_{3}$ are inside the unit circle, the Cauchy’s residue theorem gives:
$I_{4}^{(1)}(k)=2\pi i\left(\textrm{Res}[f_{4};z_{1}]+\textrm{Res}[f_{4};z_{3}]\right).$ (57)
Since both $z_{1}$ and $z_{3}$ are simple poles, their residues are easy to calculate:
$\textrm{Res}[f_{4};z_{1}]+\textrm{Res}[f_{4};z_{3}]=\frac{-i}{w_{+}-w_{-}}\left(\frac{1}{z_{1}-z_{2}}+\frac{1}{z_{4}-z_{3}}\right).$ (58)
Inserting $w_{+}-w_{-}=\frac{\sqrt{1-t_{\textrm{AM}}}}{t_{\textrm{AM}}k}\sqrt{C(k)}$, $z_{1}-z_{2}=-\sqrt{w_{+}^{2}-4}$ and $z_{4}-z_{3}=-\sqrt{w_{-}^{2}-4}$, we have
$I_{4}^{(1)}(k)=-\frac{2\sqrt{2}\pi t_{\textrm{AM}}^{2}k^{2}}{\sqrt{1-t_{\textrm{AM}}^{2}}\sqrt{C(k)}}\left[\frac{1}{\sqrt{D_{+}(k)}}+\frac{1}{\sqrt{D_{-}(k)}}\right].$ (59)
Then we evaluate $I_{4}^{(2)}(k)$ with poles shown in Fig. 3(d). To utilize the residue theorem, we construct a closed contour shown in Fig. 3(d), where the two poles $z_{1}$ and $z_{2}$ are removed by the arcs $\mathcal{C}_{r1}$ and $\mathcal{C}_{r2}$ whose radius approaches 0. To estimate the contributions from the two arcs:
$\begin{split}\lim_{z\rightarrow z_{1}}(z-z_{1})f_{4}(z)&=\frac{-i}{(z_{1}-z_{2})(w_{+}-w_{-})},\\ \lim_{z\rightarrow z_{2}}(z-z_{2})f_{4}(z)&=\frac{-i}{(z_{2}-z_{1})(w_{+}-w_{-})}.\end{split}$ (60)
According to Jordan’s lemma, the contributions of the two arcs cancel each other. Therefore:
$\begin{split}I_{4}^{(2)}(k)&=2\pi i\textrm{Res}[f_{4};z_{3}],\\ &=-\frac{2\sqrt{2}\pi t_{\textrm{AM}}^{2}k^{2}}{\sqrt{1-t_{\textrm{AM}}^{2}}\sqrt{C(k)}}\frac{1}{\sqrt{D_{+}(k)}}.\end{split}$ (61)
For all poles on the unit circle, a closed contour can be constructed in Fig. 3(e) with no singularity inside the closed contour. For the integration over the four arcs, it is straightforward to prove that the contributions from $z_{1}$ and $z_{2}$ are cancelled, so are $z_{3}$ and $z_{4}$. Therefore, $I_{4}^{(3)}(k)=0$.
V.4 Integration over $k$
After integrating the angular component $\phi$ via contour integration, our next task is to integrate the radial component $k$:
$I_{4}=-\frac{1}{(2\pi)^{2}}\frac{ab}{2}\int_{0}^{1}\frac{I_{4}(k)}{k}dk,$ (62)
where $I_{4}(k)$ can take the form $I_{4}^{(i)}(k)$ ($i=1,2,3$) according to $(k,Q)$.
Figure 4: $\frac{S(t_{\textrm{AM}},Q,0)}{\mathcal{A}}$ for $\varepsilon_{B}=0.05$ and $\mu=1.00$.
For a given $Q$, it can intersect $Q_{c1}(k)$ and $Q_{c2}(k)$ with either two or one crossings. Therefore, there are two different regimes for $Q$: (i) $Q<1/\sqrt{1-t_{\textrm{AM}}}-1/\sqrt{1+t_{\textrm{AM}}}$; and (ii) $1/\sqrt{1-t_{\textrm{AM}}}-1/\sqrt{1+t_{\textrm{AM}}}<Q<1/\sqrt{1-t_{\textrm{AM}}}+1/\sqrt{1+t_{\textrm{AM}}}$.
For $Q$ belongs to regime (i), $Q$ intersects $Q_{c1}(k)$ with two points $k_{\pm}$ whose square are given by:
$k_{\pm}^{2}=1-t_{\textrm{AM}}+\frac{t_{\textrm{AM}}}{2}(1-t_{\textrm{AM}}^{2})Q^{2}\pm\frac{1-t_{\textrm{AM}}^{2}}{2}\sqrt{4Q^{2}-(1-t_{\textrm{AM}}^{2})Q^{4}}.$ (63)
So $I_{4}$ is calculated as,
$I_{4}=-\frac{1}{(2\pi)^{2}}\frac{ab}{2}\left[\int_{0}^{k_{-}}\frac{I_{4}^{(1)}(k)}{k}dk+\int_{k_{-}}^{k_{+}}\frac{I_{4}^{(2)}(k)}{k}dk\right]=-\frac{\sqrt{2}t_{\textrm{AM}}}{2\pi}\left[\int_{0}^{k_{-}}\frac{kdk}{\sqrt{C(k)}\sqrt{D_{-}(k)}}+\int_{0}^{k_{+}}\frac{kdk}{\sqrt{C(k)}\sqrt{D_{+}(k)}}\right].$ (64)
where nest square root exists in $\sqrt{D_{\pm}(k)}$. Thanks to the additional $\sqrt{C(k)}$, such a nest square root integration can be evaluated in a closed form ($0\leq n<m\leq 1$)
$\int_{n}^{m}\frac{kdk}{\sqrt{C(k)}\sqrt{D_{\pm}(k)}}=-\frac{\sqrt{2}}{8t_{\textrm{AM}}}\ln\left|\frac{\sqrt{C(m)}\pm(1-t_{\textrm{AM}}^{2})Q+\sqrt{2D_{\pm}(m)}}{\sqrt{C(n)}\pm(1-t_{\textrm{AM}}^{2})Q+\sqrt{2D_{\pm}(n)}}\right|.$ (65)
Inserting Eq. (65) into Eq. (64), we obtain
$I_{4}=\frac{1}{8\pi}\ln\left|\frac{a-b}{a+b}\right|.$ (66)
For $Q$ belongs to regime (ii), $Q$ intersects $Q_{c1}(k)$ and $Q_{c2}(k)$ at $k_{-}$. The calculation is straightforward:
$I_{4}=-\frac{\sqrt{2}t_{\textrm{AM}}}{2\pi}\left[\int_{0}^{k_{-}}\frac{kdk}{\sqrt{C(k)}\sqrt{D_{-}(k)}}+\int_{0}^{1}\frac{kdk}{\sqrt{C(k)}\sqrt{D_{+}(k)}}\right]=\frac{1}{8\pi}\left[\ln\left|\frac{a-b}{a+b}\right|+\ln\left|\frac{\sqrt{C(1)}+(1-t_{\textrm{AM}}^{2})+\sqrt{2D_{+}(1)}}{2t_{\textrm{AM}}\sqrt{4-(1-t_{\textrm{AM}}^{2})Q^{2}}}\right|\right]$ (67)
Eq. (66) and Eq. (67) constitute the main results of this work. The most import feature of Eq. (66) and Eq. (67) is that they are invariant under $t_{\textrm{AM}}\leftrightarrow-t_{\textrm{AM}}$. Due to the duality between $I_{4}$ and $I_{3}$, such a feature gives $I_{3}=I_{4}$.
V.5 Fixed chemical potential
With $I_{2}$, $I_{3}$, and $I_{4}$ evaluated, $S(t_{\textrm{AM}},{\bf q},0)/\mathcal{A}$ now is calculated as:
$\frac{S(t_{\textrm{AM}},Q,0)}{\mathcal{A}}=\frac{1}{4\pi}\begin{cases}\ln\sqrt{\frac{2\varepsilon_{B}}{\mu}}-\ln\left|\frac{a-b}{a+b}\right|-\ln\left|\sqrt{4-(1-t_{\textrm{AM}}^{2})Q^{2}}\right|,Q\in(0,\sqrt{\frac{1}{1-t_{\textrm{AM}}}}-\sqrt{\frac{1}{1+t_{\textrm{AM}}}}]\\ \ln\sqrt{\frac{2\varepsilon_{B}}{\mu}}-\ln\left|\frac{a-b}{a+b}\right|+\ln\left|\frac{2t_{\textrm{AM}}}{\sqrt{C(1)}+(1-t_{\textrm{AM}}^{2})+\sqrt{2D_{+}(1)}}\right|,Q\in(\sqrt{\frac{1}{1-t_{\textrm{AM}}}}-\sqrt{\frac{1}{1+t_{\textrm{AM}}}},\sqrt{\frac{1}{1-t_{\textrm{AM}}}}+\sqrt{\frac{1}{1+t_{\textrm{AM}}}}]\end{cases}$ (68)
Fig. 4 plot $S(t_{\textrm{AM}},Q,0)/\mathcal{A}$ for $\varepsilon_{B}=0.05$ and $\mu=1.00$. It is clear that $S(t_{\textrm{AM}},Q,0)$ increases slowly for $Q$ in regime (i) and decreases rapidly in regime (ii), therefore $Q_{c}\equiv\sqrt{\frac{1}{1-t_{\textrm{AM}}}}-\sqrt{\frac{1}{1+t_{\textrm{AM}}}}$ is a kink for $S(t_{\textrm{AM}},Q,0)/\mathcal{A}$ which reaches the maximum value:
$\left[\frac{S(t_{\textrm{AM}},Q,0)}{\mathcal{A}}\right]_{\max}=\frac{1}{4\pi}\ln\left|\frac{\sqrt{2\varepsilon_{B}/\mu}}{\sqrt{1+t_{\textrm{AM}}}-\sqrt{1-t_{\textrm{AM}}}}\right|.$ (69)
Figure 5: (a) Fermi surface nesting for $\mathbf{q}$ along (1, 1) direction; (b) BFS for $\mathbf{q}=(q_{c}/\sqrt{2},q_{c}/\sqrt{2})$. (c) Fermi surface nesting for $\mathbf{q}$ along (1, 0) direction. (d) BFS for $\mathbf{q}=(q_{c},0)$. Here $\Delta=0$, $\mu=1$, and $t_{AM}=0.5$ for BFS plotting.
It is well known that the N phase is stable at $S(t_{\textrm{AM}},Q,0)/\mathcal{A}<0$, therefore we see that the FFLO phase emerges at $Q=Q_{c}$ and $S(t_{\textrm{AM}},Q,0)/\mathcal{A}=0$, the simultaneous solution of which leads to the critical $t_{\textrm{AM}}$ for FFLO phase:
$t_{\textrm{AM, FFLO}}=\sqrt{1-\left(1-\frac{\varepsilon_{B}}{\mu}\right){}^{2}},\frac{\mu}{\varepsilon_{B}}\geq 1$ (70)
In Fig. 2(a), we plot the FFLO phase boundary $t_{\textrm{AM, FFLO}}$ as the red curve. When $\mu<\varepsilon_{B}$, $S(t_{\textrm{AM}},Q_{c},0)=0$ has no solution and the system remains in the N phase. For $\mu\geq\varepsilon_{B}$ the FFLO phase emerges from the N phase side of the phase diagram, in agreement with our assumption that the decrease of $t_{\textrm{AM}}$ leads to a continuous N-FFLO phase transition. From Fig. 2(a), $t_{\textrm{AM}}$ has the largest window for observing FFLO phase when $\mu=\varepsilon_{B}$. After that, the window gradually shrinks.
In Soto-Garrido et al. (Soto-Garrido2014, ), the integration over $k$ is carried out within a shell around the Fermi surface, and such a finite range of integration leads to a flat $S(t_{\textrm{AM}},Q,0)/\mathcal{A}$. Based on such a result, Soto-Garrido et al. (Soto-Garrido2014, ) claimed the absence of FFLO phase. Our results reveal the importance of full $\mathbf{k}$-integration in restoring the FFLO phase.
V.6 Fixed total particle number
The upper red curve, in Fig. 2(b), is the location of the continuous phase transition at a fixed total particle number. To achieve this, we utilize $\mu=\sqrt{1-t_{\textrm{AM}}^{2}}\varepsilon_{F}$ again to obtain
$t_{\textrm{AM, FFLO}}=\sqrt{\frac{1}{2}+\frac{\varepsilon_{B}}{\varepsilon_{F}}-\sqrt{\frac{1}{4}-\frac{\varepsilon_{B}}{\varepsilon_{F}}}},\frac{\varepsilon_{B}}{\varepsilon_{F}}\leq\frac{1}{4}$ (71)
This is shown as the upper red curve in Fig. 2(b). It is clear that the larger $\varepsilon_{B}/\varepsilon_{F}$ is, the wider the $t_{\textrm{AM}}$ window for observing FFLO phase. Such a feature greatly facilitates the preparation of FFLO states in electronic superconductors.
V.7 Geometric interpretation
In previous section, we have shown that, for $\Delta=0$, $S(t_{\textrm{AM}},Q,0)/\mathcal{A}$ exhibits a nonanalytic dependence on $Q$ according to the contour integration of complex analysis. The nonanalyticity occurs at $Q_{c}=\sqrt{\frac{1}{1-t_{\textrm{AM}}}}-\sqrt{\frac{1}{1+t_{\textrm{AM}}}}$ which corresponds to $q_{c}=a-b$. As shown in Fig. 5(a), such a $q_{c}$ is the critical momentum that two of the four crossings between the two Fermi ellipses ($E_{{\bf k}\pm}=A({\bf k})\pm B({\bf k})=0$) and the black dashed circle ($A({\bf k})=0$) becomes one tangent point.
We note that the black dashed circle in Fig. 5(a) is where the denominator of the first term in Eq. (44) vanishes. i.e. it represents gapless excitations. Therefore the onset of FFLO phase here is accompanied by the decrease of gapless excitations. Such a fact is in huge contrast with FFLO phase driven by magnetic field, where FFLO phase emerges when the crossing point increase from 0 to 1.
The physical meaning of $q_{c}$ can also be understood through the robustness of the BFS. Since the BFS is protected by the topological invariant $l$, it remains stable until a topological phase transition occurs. This transition is precisely what happens at $q_{c}$: as shown in Fig. 5(b), the number of BFS decreases from 4 to 3 at this critical momentum.
Motivated by the above analysis, we may determine $q_{c}$ when $\mathbf{q}$ is along the $(1,0)$ (or equivalently $(0,1)$) direction, where obtaining a closed form of $S(t_{\textrm{AM}},Q,0)/\mathcal{A}$ is difficult. Fig. 5(c) demonstrates the condition when the number of crossings decreases from 4 to 2, and the corresponding $q_{c}$ is
$q_{c}=\frac{\sqrt{2}(a^{2}-b^{2})}{\sqrt{a^{2}+b^{2}}}=2\sqrt{\frac{\mu t_{\mathrm{AM}}^{2}}{1-t_{\mathrm{AM}}^{2}}}.$ (72)
As elaborated in Appendix A, our numerical simulation verifies that Eq. (72) indeed gives a good estimate of the actual $q_{c}$ when $\mathbf{q}$ is along the $(1,0)$ direction.
VI CONCLUSIONS AND OUTLOOKS
In summary, we have analytically investigated the ground-state phase diagram of 2D spin-1/2 Fermi gases exhibiting $d$-wave altermagnetic spin splitting in the presence of $s$-wave pairing, within the framework of BdG mean-field theory. Under both fixed chemical potential and fixed total particle number conditions, we identify four distinct phases as illustrated in Fig. 2. Our results definitely confirm the existence of FFLO states under $s$-wave pairing, thereby resolving previous contradictory conclusions in the literatures. In addition, we provide a geometric interpretation of the onset $\mathbf{q}_{c}$ of FFLO states, offering intuitive insight into its origin. Quite unexpectedly, we also discover a nodal SF, characterized by the TBFS. Given the simplicity of the system - a single band, without external magnetic field, and under nodeless $s$-wave pairing - the presence of TBFS is both surprising and compelling.
Fig. 2 illustrates the fundamental relationship between anisotropic NRSS and attractive pairing. When the attractive pairing is strong, BCS state is expected with a zero-momentum and fully gapped Fermi surface. Meanwhile, for a very large anisotropic NRSS, the metallic normal phase is anticipated. Nodal SF and FFLO occur when anisotropic NRSS and attractive pairing are both at intermediate strength. All phases can be realized in either ultracold gases or electronic superconductors by choosing an appropriate $t_{\textrm{AM}}$. In the context of ultracold gases, a large $t_{\textrm{AM}}$ is required as shown in Fig. 2(a), so we can access to all four phases via tuning the two-body interactions $\varepsilon_{B}$ through Feshbach resonance Chin2010 . For electronic superconductors, a small $t_{\textrm{AM}}$ suffices, as shown in Fig. 2(b). By increasing the Fermi energy $\varepsilon_{F}$ via techniques like electrostatic gating Wu2023 , we can sequentially observe BCS, nodal SF, FFLO, and normal phases.
We conclude this section by noting several natural extensions of this work. First, incorporating additional physical ingredients - such as external magnetic fields, spin-orbit coupling, $d$-wave pairing, and finite temperature - could significantly enrich the phase diagram and reveal new emergent pairing phenomena. Second, while our analysis is based on the simplest single-plane-wave ansatz here, higher plane-wave FFLO ansatz such as the LO state with $\Delta(\mathbf{x})=\Delta\cos(\mathbf{q}\cdot\mathbf{x})$ are expected to be energetically more favorable. Finally, the topological BFS uncovered in this work may provide a fertile platform for realizing Majorana bound states, opening the door to potential applications in topological quantum computation.
Acknowledgements.
We thank Jia Wang for helpful discussions. This research was supported by the Australian Research Council’s (ARC) Discovery Program, Grants Nos. DP240101590 (H.H.) and DP240100248 (X.-J.L.).
Appendix A Numerical calculations of the inverse vertex function $S(t_{\textrm{AM}},\mathbf{q},\omega=0)$
In the inverse vertex function $S(t_{\textrm{AM}},\mathbf{q},\omega=0)$, the two-dimensional integration over the momentum $\mathbf{k}$ could be performed numerically. In Fig. 6, we present a comparison between the numerical results and the analytic expression given in Eq. (68), for the case where the center-of-mass momentum $\mathbf{q}$ lies along the diagonal direction (i.e., the $(1,1)$ direction with $\phi=\pi/4$), at the binding energy $\varepsilon_{B}=0.08\varepsilon_{F}$. On the scale of the figure, the two theoretical predictions appear nearly indistinguishable, even though a small temperature $T=0.005\varepsilon_{F}/k_{B}$ has been introduced in the numerical calculation to smooth out the sharpness of the Fermi surface.
Figure 6: The inverse vertex function $S(t_{\textrm{AM}},q,\omega=0)$ as a function of the center-of-mass momentum $q$, along the $x$-axis $(1,0)$ direction (red dot-dashed line) and along the diagonal $(1,1)$ direction (black solid line and blue circles). Here, we consider an altermagnetic coupling strength $t_{\textrm{AM}}=0.41$. The numerical calculations are carried out at a negligible temperature $T=0.005\varepsilon_{F}/k_{B}$, with a binding energy $\varepsilon_{B}=0.08\varepsilon_{F}$ and a chemical potential $\mu=\sqrt{1-t_{\textrm{AM}}^{2}}\varepsilon_{F}\simeq 0.912\varepsilon_{F}$.
Figure 7: A three-dimensional plot of the inverse vertex function $S(t_{\textrm{AM}},\mathbf{q},\omega=0)$ as functions of $q_{x}$ and $q_{y}$. The parameters used in the numerical calculations are the same as in Fig. 6.
In Fig. 6, we also present the numerical result for the case where the center-of-mass momentum $\mathbf{q}$ is along the $x$-direction with $\phi=0$, a configuration for which an analytic expression is not available. Interestingly, we observe that the inverse vertex function reaches its maximum at $\phi=0$, rather than at $\phi=\pi/4$. This trend is most clearly illustrated in the three-dimensional plot of the inverse vertex function shown in Fig. 7. Despite this angular dependence, the peak values of $S(t_{\textrm{AM}},\mathbf{q},\omega=0)$ at different $\phi$ are relatively close, implying that the critical altermagnetic coupling strength $t_{\textrm{AM,FFLO}}$ in Eq. (70) and Eq. (71) remains highly accurate. To verify this, we numerically determine $t_{\textrm{AM,FFLO}}(\phi=\pi/4)=0.409$, which is in excellent agreement with the analytic result from Eq. (71), and $t_{\textrm{AM,FFLO}}(\phi=0)=0.433$. The small discrepancy between these two critical values confirms that it is reasonable to approximate
$t_{\textrm{AM,FFLO}}$ $=t_{\textrm{AM,FFLO}}(\phi=0),$ (73)
$\simeq t_{\textrm{AM,FFLO}}(\phi=\pi/4).$ (74)
Therefore, the red curves in the phase diagram Fig. 2 are nearly unchanged, if we take into account the angular dependence of $S(t_{\textrm{AM}},\mathbf{q},\omega=0)$ on the momentum $\mathbf{q}$.
Finally, we check the critical momentum $q_{c}$ when the center-of-mass momentum $\mathbf{q}$ is oriented along the $(1,0)$ direction ($\phi=0$). From the red dashed line in Fig. 6, we estimate $q_{c}(\phi=0)\simeq 0.80k_{F}$. This numerically determined value is in good agreement with the analytic prediction Eq. (72) from the geometric consideration, which yields $q_{c}(\phi=0)\simeq 0.86k_{F}$ at $t_{\textrm{AM}}=0.41$.